Bond valuation—Semiannual interest Find the value of a bond maturing in 9 years, with a $1,000 par value and a coupon interest rate of 14% (7% paid semiannually) if the required return on similar-risk bonds is 14% annual interest (7% paid semiannually). 


The present value of the bond is $_______ 
Answer #1

Cash Inflow Per Half Year = $ 1000 * 7% = $70

Rate of PVIF = 14/2 = 7% ( because compounding semiannually)

total no of years = 18 because compounding Semiannually

Year Cash Inflow PVIF 7% Present Value of Cash Inflow
(A) (B) C= A*B
1 $70 0.9346 65.42
2 $70 0.8734 61.14
3 $70 0.8163 57.14
4 $70 0.7629 53.40
5 $70 0.7130 49.91
6 $70 0.6663 46.64
7 $70 0.6227 43.59
8 $70 0.5820 40.74
9 $70 0.5439 38.08
10 $70 0.5083 35.58
11 $70 0.4751 33.26
12 $70 0.4440 31.08
13 $70 0.4150 29.05
14 $70 0.3878 27.15
15 $70 0.3624 25.37
16 $70 0.3387 23.71
17 $70 0.3166 22.16
18 $70 0.2959 20.71
18 $1,000 0.2959 295.90
TOTAL $1,000

Value of Bond = $1000 proved

Alternatively Ans is

When Risk free rate and Interest/coupon Rate is Same then Value of Bond is always equal to its Face Value , In this Case ie $ 1000